OpenStudy (anonymous):

Given that sin 2pi/7 = 0.7818, use equivalent trigonometric expressions to evaluate cos 3 pi/14

OpenStudy (anonymous):

cos 3pi/14 = 0.7818 \[\cos (\pi/2 - 3\pi/14) = \cos2\pi/7\] since sin and cosine give same angles, sin 2pi/7 is the same as cos 2pi/7.